OpenStudy (janu16):

how to find zeros of –x4 + x3 + 7x2 − x − 6

OpenStudy (kenshin):

trial and error for a root first, then use long division once you've found a root

OpenStudy (janu16):

but i dont know how to do tht

OpenStudy (kenshin):

make -x^4+x^3 +7x^2 - 6 = 0, now trial and error numbers as x value see which one when subbed in, gives you 0.

OpenStudy (kenshin):

oops i meant your equation -x^4 + x^3 + 7x^2 - x - 6 = 0

OpenStudy (janu16):

1 works

OpenStudy (janu16):

i used synthetic division to find that out

OpenStudy (janu16):

i mean -1

OpenStudy (kenshin):

yes, so when x = 1, the entire equation = 0, so you could say x - 1 = 0 is a root of the equation, since the equation has a total of power 4, you can divide the equation by (x-1) and get a power 3 equation

OpenStudy (kenshin):

ok then it's x+1 = 0

OpenStudy (kenshin):

so (x+1) (ax^3+bx^2+cx+d) = 0

OpenStudy (kenshin):

when your -1 works, you gotta move it to the other side of the equation to change its sign. for instance when x = -1 the whole equation = 0, moving -1 to the left by adding 1 on both sides give you x+1 = 0

OpenStudy (janu16):

ohok s

OpenStudy (janu16):

so what do you do next

OpenStudy (kenshin):

you can use your synthetic division, or it's easier to just use long division to divide -x^4+x^3+7x^2 - x -6 by (x+1), and get your cubic equation

OpenStudy (janu16):

how would you solve tat with long division

jhonyy9 (jhonyy9):

@Kenshin --- ATTENTION please ! above you have wrote this :,,yes, so when x = 1, the entire equation = 0, so you could say x - 1 = 0 is a root of the equation, since the equation has a total of power 4, you can divide the equation by (x-1) and get a power 3 equation - so than how is true ? divide by x-1 or x+1 ?

OpenStudy (kenshin):

in this case both works, so there are 2 roots of -1 and +1, so you could in fact divide the whole equation by x^2-1 and get a much simpler quadratic quotient as a result.

jhonyy9 (jhonyy9):

so than you try x=-1 this is sure that you will get = zero ?

OpenStudy (janu16):

so i really need with this like can you write in like step after i did my synthetic divison cause i dont get it and i need help

jhonyy9 (jhonyy9):

so i think in case of -1 you will get 2=0

jhonyy9 (jhonyy9):

so what is false - or not ?

jhonyy9 (jhonyy9):

@Kenshin - are you here ?

OpenStudy (kenshin):

-(-1^4) +(-1^3) +7(-1^2) - (-1) - 6 -1-1+7+1-6=0

OpenStudy (kenshin):

That video teaches you how to do long division with algebra Janu16

OpenStudy (janu16):

ok

OpenStudy (janu16):

so can i get my 4 roots by doing long division

OpenStudy (kenshin):

since both -1 and +1 works as roots, you will get a quadratic (ax^2+bx+c)*(x+1)*(x-1) once you divide the whole equation by x^2 - 1 yes

OpenStudy (kenshin):

but whether it's 4 roots or 3 roots we don't know that yet, it could be a perfect square that has the curve touching a point instead of crossing it as well in that case there's a double root and those 2 roots which makes them 3 roots.

OpenStudy (kenshin):

ok do you see how it works now? can you complete the rest? 

OpenStudy (janu16):

yea im working

OpenStudy (kenshin):



OpenStudy (janu16):

never learned long polynomial so having hard time so do you do x^2 times x^3?

OpenStudy (kenshin):

no, i completed the first step for you which was: x^2 times (x^2 - 1) [that was your 2 roots you got from (x+1)(x-1)]; and you get x^4 - x^2 to take away from your original equation, then you gotta find the next term at the top to time by (x^2 - 1) to try and cancel out the remainder, repeat this process until you reach 0

OpenStudy (kenshin):

sorry -x^2 i meant